A car is traveling at 60.0 km/hr. The wheels on the car are rolling without slipping along level ground. The radius of each wheel is 0.5 m. What is the angular velocity, ω, of each wheel, since they are all the same size and are traveling at the same forward speed?
Answer #1

As they are rolling without sleeping we can use linear relation to express velocity and angular velocity

linear velocity is

v=60\ km/hr=\frac{60\times1000}{60\times60}=16.67\ m/s

Angular velocity corresponding to a linear velocity v of an object of radius r is

\omega=\frac{v}{r}=\frac{16.67}{0.5}\ rad/s=33.34\ rad/s